Describe at least three ways that you can maximize the growth of your investment over time. Do the following calculations to help you determine what factors will help your money to grow, and use your results from these calculations in your response to this question.

  • invest $5000 at 4.5% simple interest versus 4.5% compound interest (quarterly), and compare the results after 5 years
  • invest $2500 at 4.5% compound interest versus $5000 at 4.5% compound interest, and compare results after 5 years with quarterly compounding
  • invest $5000 at 4.5% compounded monthly versus 8% compounded monthly, and compare results after 5 years
  • invest $5000 at 4.5% compounded monthly for 5 years versus 10 years, and compare results

Again, the purpose of this question is NOT for you to simply do calculations, but to analyze your results, and explain what you can do as an investor to help your money grow. Your response should be a paragraph with numbers in it, not simply numbers.


1
Expert's answer
2021-10-21T09:56:20-0400

1.

simple interest:

"FV=5000(1+0.045)^5=\$6230.91"

quarterly interest:

"FV=5000(1+0.045/4)^{5\cdot4}=\$6253.75"